OpenStudy (anonymous):

Verify the identity. Show your work. 1 + sec2xsin2x = sec2x

OpenStudy (ronlover101):

I like your pic lucy is it you???

OpenStudy (solomonzelman):

\(\large\color{black}{ \displaystyle 1 + \sec^2x\sin^2x = \sec^2x }\) Like this?

OpenStudy (solomonzelman):

I would multiply everything times \(\color{black}{\cos^2x}\). That would give me the pathegorean identity. But, if you are only allowed to play with one side at a time, I would factor out of \(\color{black}{\sec^2x}\) on the left side.

OpenStudy (solomonzelman):

Hint: \(\Large\color{black}{ \displaystyle \color{red}{1} + \sec^2x\sin^2x = \sec^2x }\) \(\Large\color{black}{ \displaystyle \color{red}{\sec^2x\cos^2x} + \sec^2x\sin^2x = \sec^2x }\) Look, \(\Large\color{black}{ \displaystyle \color{blue}{\underline{\color{black}{\sec^2x}}}\cos^2x +\color{blue}{\underline{\color{black}{\sec^2x}}}\sin^2x = \color{green}{\underline{\color{black}{\sec^2x}}} }\)

OpenStudy (anonymous):

i just dont get it